OpenStudy (anonymous):

The graph below has the same shape as the graph of G(x) = x2, but it is shifted down three units and to the left two units. Complete its equation. Enter exponents using the caret (^); for example, enter x2 as x^2. Do not include "F(x) =" in your answer.

OpenStudy (anonymous):

i tryed to do it on my own but didnt work and looked at examples and still didnt work.

OpenStudy (anonymous):

hello ?

OpenStudy (anonymous):

g(x)= x2 is an upward parabola with axis (0,0)... have u got that??

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

if u put Y= y-3 ... and X= x ???

OpenStudy (anonymous):

have you seen the picture

OpenStudy (anonymous):

I could not see it...access denied...sad times

OpenStudy (anonymous):

thtss weird it let me see it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You can attach the image to the problem directly rather than linking it.

OpenStudy (anonymous):

Ok, so the graph is the same as \(y = x^2\) But it's shifted down 3 units. So how will the y values of the graph relate to the y values of \(y=x^2\) for a particular value of x?

OpenStudy (anonymous):

can i just have the answer just to see what you put tht i didnt put i had (x+-3)^2

OpenStudy (anonymous):

If we call F(x) the y value of the function and G(x) = \(x^2\). What is F(x) in terms of G(x)?

OpenStudy (anonymous):

idk , (x+1)^2

OpenStudy (anonymous):

If x = 0, what will F(x) be?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Is that what your graph shows? Here's a better question. If H(x) is just had a parabola that looked exactly like \(x^2\) but was shifted down 3 units. What would H(0) be?

OpenStudy (anonymous):

\(0^2 = 0,\) but our H(x) is shifted down 3 from this graph, so H(0) would be?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

And if x = 1?

OpenStudy (anonymous):

What is \(G(x) = x^2 =?\) and what is our hypothetical H(x) ?

OpenStudy (anonymous):

(1+-3)^2

OpenStudy (anonymous):

No. If x=1, then \(G(x) = (1)^2 = 1\) And since H is still shifted down 3, we'd have H(x) = 1-3 = -2. So we can see that for any given value for x, H(x) = G(x) - 3

OpenStudy (anonymous):

look i dont knoww this stuff , im just about to give im seriously

OpenStudy (anonymous):

You're doing fine, but I'm trying to take it one step at a time and lead you to the concepts that will allow you to solve these quickly and easily. I think part of the problem is that you imagine that you just aren't going to be able to understand and therefore try to just guess at a solution to get it over with.

OpenStudy (anonymous):

Do you have a little bit of time?

OpenStudy (anonymous):

yess , alot

OpenStudy (anonymous):

Ok. Lets start over then. Forget about this problem for a second. And lets take a look at a simpler problem.

OpenStudy (anonymous):

If I have a function G(x) = x, What does that look like?

OpenStudy (anonymous):

yess , alot

OpenStudy (anonymous):

F(x)=x

OpenStudy (anonymous):

No, G(x) = x. What does the graph of this look like?

OpenStudy (anonymous):

I haven't said anything about F(x) yet ;) be patient =)

OpenStudy (anonymous):

look idk , its crazzy i just need the answer so i can move on i know that your trying to help me and to get better at my math but i just need the answer

OpenStudy (anonymous):

Getting the answer will not help you on the test.

OpenStudy (anonymous):

it will .

OpenStudy (anonymous):

You said you have time. Please also have patience also.

OpenStudy (anonymous):

this is in a time thing so yeaa .

OpenStudy (anonymous):

G(x) = x is just a line passing through the origin. When x = 0, G(x) = 0, when x = 1, G(x) = 1, and so on.

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

okay that was more simple next step

OpenStudy (anonymous):

Now lets say we have another function F(x) that is the same shape (a straight line with a slope of 1) but it's not going through the origin. It is shifted down 3 units. \(x = 0 \implies F(x) = ? \) \(x = 1 \implies F(x) = ? \)

OpenStudy (anonymous):

Given how I have discribed F, what will F(0) be?

OpenStudy (anonymous):

If G(0) = 0, and F(0) is the same as G(0) but shifted down 3 units.. What is F(0)?

OpenStudy (anonymous):

look im geting confused , this math is stressing mee outt , like really

OpenStudy (anonymous):

Deep breath. If G(0) = 0, and F(0) is G(0) but shifted down 3 units, wouldn't F(0) = 0 - 3 = -3?

OpenStudy (anonymous):

ive been said -3

OpenStudy (anonymous):

You said -3 about something else. I said we were starting over with an easier problem.

OpenStudy (anonymous):

What is F(1) ?

OpenStudy (anonymous):

manee , can you just tell me the answer because ive been stuck on the same question for like an hour

OpenStudy (anonymous):

If you are not interested in learning, I'm not interested in teaching. No I won't help you cheat.

OpenStudy (anonymous):

then flutter you then lame